OpenStudy (anonymous):

ln(-4-x)+ln3=ln(2-x) every time i try it i get no solution because the x's cancel out. help?

OpenStudy (anonymous):

\[\ln(-4-x) + \ln(3) = \ln(-12-3x) = \ln(2-x)\] so \[-12 - 3x = 2 - x \rightarrow 2x = -14\space \text{so}\space x= -7\]